OpenStudy (anonymous):

Josie Makes Fruit punch by mixing fruit juice and lemonade in the ratio 2:3 she needs to make 30 litres of punch for a party How much of each ingrediants does she need? Fruit Juice=..... Litres Lemonade=.....Litres During The party Josie Decides to make some more she has 6 litres of fruit juice left and plenty of lemonade How much extra punch can she make

OpenStudy (theviper):

I m going to try OK :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I Know the first bit but i dont know this bit During The party Josie Decides to make some more she has 6 litres of fruit juice left and plenty of lemonade How much extra punch can she make

OpenStudy (theviper):

According to question, ratio is 2:3 than 2x+3x=30 5x=30 x=30/5=6 Therefore, FRUIT JUICE=2x=2*6=12 liters & Lemonade=3x=3*6=18 liters

OpenStudy (theviper):

Am i right?

OpenStudy (anonymous):

Yh

OpenStudy (anonymous):

Could u help me with the second bit

OpenStudy (anonymous):

During The party Josie Decides to make some more she has 6 litres of fruit juice left and plenty of lemonade How much extra punch can she make

OpenStudy (theviper):

is the answer of second is \(15\) ?

OpenStudy (anonymous):

yup...u r right

OpenStudy (anonymous):

could you help with with the third bit To make the second batch of punch go further Josie adds 2 more litres of lemonade What is the ratio of fruit juice to lemonade now

OpenStudy (theviper):

I think 6:11

OpenStudy (theviper):

right @Yahoo! ?

OpenStudy (anonymous):

is it correct Yahoo

OpenStudy (theviper):

.com :D

OpenStudy (anonymous):

Loool

OpenStudy (theviper):

Are u here @Yahoo! ?

OpenStudy (anonymous):

innih

OpenStudy (theviper):

I think ur question is solved:S

OpenStudy (anonymous):

Yh

OpenStudy (theviper):

:)

OpenStudy (anonymous):

Can i ask another one

OpenStudy (theviper):

yup by closing this :)

OpenStudy (anonymous):

Mrs Perry shares out 15 biscuits between Gemma and Zak in the ratio 1:2 Howmany biscuits does each child get Gemma gets...... Biscuits and Zak get........ Biscuits

OpenStudy (theviper):

first plz close this question :)

OpenStudy (anonymous):

kl

OpenStudy (anonymous):

@James32841234 this question is similar to the one which u posted now..Plzz....Try Doing it by ur Own...:)

OpenStudy (anonymous):

Follow the Same Steps...:)

OpenStudy (anonymous):

@TheViper Plzzz...Make him to Do..it...:)

OpenStudy (anonymous):

No its to difficult

OpenStudy (theviper):

1x+2x=15 3x=15 x=? therfore Gemma gets=1x=? Zak gets=3x?

OpenStudy (anonymous):

Wats the answer

OpenStudy (theviper):

Solve that

OpenStudy (anonymous):

Um--um---...Do it by ur Own...

OpenStudy (anonymous):

It is Simple Algebra

OpenStudy (theviper):

Very Simple :)

OpenStudy (anonymous):

I dont know the answer PLZ

OpenStudy (theviper):

If you do not do it by ur own than you are going to face problems in other Competitions

OpenStudy (anonymous):

Plz

OpenStudy (theviper):

ok i m going to make it easy:)

OpenStudy (anonymous):

Thanks

OpenStudy (theviper):

1x+2x=15 3x=15 x=15/3=?

OpenStudy (anonymous):

Dont Know plz tell me the answer

OpenStudy (theviper):

x=5 therefore Gemma gets=1x=1*5=5 & Zak gets=2x=2*5=10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can i ask u another question

OpenStudy (anonymous):

Mr Castle Shares out 64 marbles Emily and Asif in the ratio 7:1How many Marbles does each child get? Emily gets...... Marbles and Asif Gets.....Marbles